8/y= 17/3 solve for y and rounded to the nearest tenth
Answer by checkley75(3666)

8/Y=17/3 NOW CROSS MULTIPLY THUS:
17Y=8*3
17Y=24
Y=24/17
Y=1.4 ANSWER.